OpenStudy (anonymous):

I have to find the rate at which a runner burns calories? the numbers on the graph are 11 and 18, i need to find how many calories per mintue

OpenStudy (anonymous):

I assume these are points?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

what are x values

OpenStudy (anonymous):

there is none its just give you the 2 points that are on the graph

OpenStudy (anonymous):

which are 11 and 15

OpenStudy (anonymous):

is there anyway you can scan the graph?

OpenStudy (anonymous):

i can try. give me one sec

OpenStudy (anonymous):

here is the graph, you might have to make it bigger

OpenStudy (anonymous):

your two points are (10,80) (14,120)

OpenStudy (anonymous):

can you explain how you got that please?

OpenStudy (anonymous):

i am sorry i haven't taken math in soo long since i was in highs school

OpenStudy (anonymous):

there were two dots, I looked at x and y value of each dot

OpenStudy (anonymous):

oh, ok i see what u are talking about now

OpenStudy (anonymous):

find slope between two points slope= y2 - y1 ------- x2-x1 (10,80) (14,120) slope= 120 - 80 ------- 14-10 40/4=10 so 10 calories per minutes

OpenStudy (anonymous):

where the one coming from. are we replacing the one?

OpenStudy (anonymous):

10 was not the right answer, it says its wrong

OpenStudy (anonymous):

I think its -10

OpenStudy (anonymous):

that was wrong too

OpenStudy (anonymous):

may be we need to recheck points chosen

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i will be right back

OpenStudy (anonymous):

back

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

yeah,let me relook at the graph

OpenStudy (anonymous):

i am looking at it too

OpenStudy (anonymous):

I think it might be (15,120) instead of (14,120) find slope between two points slope= y2 - y1 ------- x2-x1 (10,80) (15,120) slope= 120 - 80 ------- 15-10 40/5=8 so 8 calories per minutes

OpenStudy (anonymous):

that was right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

gotta get my eyes checked out

OpenStudy (anonymous):

yea me too lol

OpenStudy (anonymous):

how can i identify the y-intercept

OpenStudy (anonymous):

it is y value when x=0 in your graph , when x=0 , y=0 as well

OpenStudy (anonymous):

oh, ook